OpenStudy (anonymous):

each of my daughters has as many brothers as sisters, but each of my sons has three times as many sisters as brothers. how many children do i have?

OpenStudy (anonymous):

for #27, you know the lines mean they are equivalent right? so you can sum them up to get the perimeter. and the left side would just be 5mm x 3.

OpenStudy (mani_jha):

Let the no of sons and daughters be x and y. Now, each of the y daughters has (y-1) sisters and x brothers. Agree? So, y-1=x Also, each of the x sons has x-1 brothers and y sisters. So, y=3(x-1) Did you understand? Now, can you solve for x and y?

OpenStudy (mani_jha):

First of all, did you understand how I got those two equations?